# Assessing Universal Relations for Rapidly Rotating Neutron Stars: Insights from an Interpretable Deep Learning Perspective

### v1

@@ Abstract
Relations between stellar properties independent of the nuclear equation of state offer profound insights into neutron star physics and have practical applications in data analysis by breaking the degeneracy between the parameters of interest. Commonly, these relations are derived from utilizing various realistic nuclear cold hadronic, hyperonic, and hybrid EoS models, each of which should obey the current multimessenger constraints and cover a wide range of stiffnesses. Concurrently, the field of multimessenger astronomy has been significantly enhanced by the advent of gravitational wave astronomy, which increasingly incorporates deep learning techniques and algorithms. At the same time, X-ray spectral data from NICER based on known pulsars are available, and additional observations are expected from upcoming missions. Therefore, in the coming years, a wealth of information on neutron stars is expected to be gathered from multiple observational channels. In this study, we revisit established universal relations, introduce new ones, and reassess them using a feed-forward neural network as a regression model. More specifically, we mainly propose “deep” EoS-insensitive hypersurface relations for rapidly rotating compact objects between several of the star’s global parameters, such as the moment of inertia, the reciprocal of the moment of inertia, the rotational frequency, mass and current multipole moments and the equatorial radius, which achieve an accuracy of within $1\%$ in most cases, with only a small fraction of investigated models exceeding this threshold. While analytical expressions can be used to represent some of these relations, the neural network approach demonstrates superior performance, particularly in complex regions of the parameter space. Furthermore, we use the SHapley Additive exPlanations (SHAP) method to interpret the suggested feed-forward network’s predictions, since SHAP is based on a strong theoretical framework inspired by the field of cooperative Game Theory. Most importantly, these highly accurate universal relations empowered with the interpretability description could be used in efforts to constrain the high-density equation of state in neutron stars, with the potential to enhance our understanding as new observables emerge.
@@ III.1 Mapping explainability using the SHAP method
Unfortunately, recent ML problems are becoming increasingly challenging and require more advanced and complex models to tackle them [184]. However, model complexity typically presents a trade-off between explainability and performance in decision-making. For instance, simpler models like those employing linear regression are highly interpretable but typically fall short in predictive performance compared to DL ones. In addition, DL models often sacrifice interpretability due to their inherent complexity, and nonlinear transformations, among other factors [185]. On top of that, contrary to the common perception that DL models operate as “black-boxes”, they are, in fact, well-defined systems, and we can provide clear explanations for how the input parameters influence their behavior. It should be noted that in several cases, understanding why a model made a certain prediction can be crucial.
Over the years, various techniques have been developed to interpret the predictions of complex models, including LIME, DeepLIFT, and SHAP [154]. Among these, SHapley Additive exPlanations (SHAP) is a widely recognized method inspired by the concept of Shapley values from the field of cooperative game theory [155, 156]. Building on this foundation, it is essential to demonstrate how Shapley values, denoted as $\phi_{i}$, provide a systematic quantification of the contribution of each feature (model’s input parameter) to the model’s predictions. Following this approach, we consider a set of features $D$ and a function $\mathcal{F}$ that maps subsets of these features to real numbers, defined as $\mathcal{F}:2^{D}\rightarrow\mathbb{R}$, with $\mathcal{F}(\emptyset)=0$, where $\emptyset$ represents the empty set. The contribution of a specific feature $f_{i}$ to the mapping $\mathcal{F}$ is then quantified by its Shapley value $\phi_{i}$, defined as:
$\phi_{i}(D,\mathcal{F})=\sum_{S\subseteq D\setminus\{f_{i}\}}\frac{|S|!(d-|S|-1)!}{d!}(\mathcal{F}(S\cup\{f_{i}\})-\mathcal{F}(S)),$ (4)
where $d$ denotes the total number of features. The summation iterates over all subsets $S$ of $D$ that exclude $f_{i}$, including the empty set. In essence, computing each Shapley value requires evaluating all possible feature subset combinations that exclude the specific feature $f_{i}$. For each subset, the contribution of $f_{i}$ is determined by the weighted marginalization term ($\mathcal{F}(S\cup\{f_{i}\})-\mathcal{F}(S)$), as defined in Eq. (4). This approach provides a rigorous framework for assessing the contribution of each feature to the ML model’s predictions.
@@
$\hat{\mathcal{F}}(x)=\mathbb{E}_{X}(\hat{\mathcal{F}}(X))+\sum_{j=1}^{D}\phi_{j},$ (5)
where $\hat{\mathcal{F}}(x)$ represents the model’s prediction, $\mathbb{E}_{X}(\hat{\mathcal{F}}(X))$ is the model’s expected value, and $\phi_{j}$ are the individual SHAP values. In Fig. 1, as an indicative clarification example, we present the contribution of each SHAP value $\phi_{i}$ regarding the specific input vector $x_{i}=(f_{i,1},f_{i,2},f_{i,3},f_{i,4})$, containing four features. Each SHAP value contributes positively (represented by a green-colored vector) or negatively (represented by a red-colored vector) to the model’s inference, $\hat{\mathcal{F}}(x_{i})$, relatively to $\mathbb{E}_{X}(\hat{\mathcal{F}}(X))$. These values serve as useful tools to explain the model’s prediction ability, offering interpretability to the deep learning model’s predictions.
@@
In summary, interpreting the results of a deep learning (DL) model using SHAP values provides valuable insights into the factors influencing its predictions. SHAP values offer a systematic approach to quantifying the contribution of each input feature to the model’s output, thereby aiding in understanding its decision-making process. By analyzing SHAP values, we can identify which features have the most significant impact (in terms of magnitude) and how they contribute (positively or negatively) relative to the model’s expected value. This interpretability framework not only enhances our understanding of the model’s behavior but also supports its validation and helps identify potential areas for improvement or further investigation.
@@ III.2 Deep Learning Architecture, Training and Testing
$\mathcal{L}_{\delta}(z,\hat{z}(\theta))=\begin{cases}\frac{1}{2}\sum_{i=1}^{n}|z_{i}-\hat{z}_{i}(\theta)|^{2}&\text{, }|z_{i}-\hat{z}_{i}|\leq\delta,\\ \\ \delta\left(\sum_{i=1}^{n}|z_{i}-\hat{z}_{i}(\theta)|-\frac{1}{2}\delta\right)&\text{, }|z_{i}-\hat{z}_{i}|>\delta,\end{cases}$ (6)
@@
For designing a DL architecture that satisfies our requirements, we utilized the PyTorch library [189]. For each EoS of cold, dense nuclear matter included in our catalog, we have used a random selection of 25% rotating stellar models as a test set333random.seed(42) method implemented in Python was used to reproduce the same sample of random NS configurations for testing. This particular choice ensures that the results are consistent across different runs.. In our systematic investigation, we employed an ANN architecture described by an input layer encompassing the features [^4] Depending the case, the specific characteristics of these features will be clarified in the Sec. IV. $x=(f_{1},...,f_{D})$, followed by five hidden layers denoted as $H_{1},\ldots,H_{5}$, while concluding in a singular output layer represented by $\hat{z}$. The number of the associated neurons regarding each hidden layer is shown in Table (1).
@@
to map each input feature $x_{i}$ to the interval $[0,1]$. This process is carried out on the entire dataset, which includes the associated quantities. Feature scaling is essential to ensure that all features contribute equally, preventing the dominance of those with larger values. This practice enhances the convergence and overall performance of the algorithms used. Either way, this process is an important step to render the learning process faster. Depending on the universal relation, we employ an adaptive learning rate, a tailored learning rate strategy, a case-favored mini-batch size, and a specific number of epochs for each training phase. In each case, the training time is linearly proportional to the number of epochs.
Across all of the regression models examined, we employed comprehensive evaluation measures to assess the performance of each ANN across the associated test sets (see e.g., [143, 153] for a detailed review). The statistical measures utilized include Mean Absolute Error (MAE), Max Residual Error (MRE), Mean Squared Error (MSE), Max Relative Deviation ($\mathrm{d}_{\text{max}}$), Mean Absolute Percentage Error (MAPE), Explained Variance, and Determination Coefficient ($R^{2}$) [^5] For more details about the statistical evaluation functions the interested reader can be directed to this link: https://scikit-learn.org/stable/modules/model_evaluation.html. Notably, a marginal increase in evaluation measures was observed in the validation set compared to the training set. In each case, as is standard practice, we report the evaluation measures based on the performance observed on the test datasets.
@@
$\bar{M_{4}}$ 2.655 497.192 2.658 491.585
$\bar{M_{4}}/\bar{S}_{3}$ 1.525 16.021 1.524 15.961
@@
Last but not least, we define a relation between the NS parameters as EoS-insensitive when the selected features at the DL model’s inference yield relative errors in the training and test sets of $\lesssim\mathcal{O}(10\%)$. Crucially, our regression “deep” models accurately reproduce the data on the test set, showcasing a state-of-the-art generalization capability beyond the training data. For each universal relation examined, the ANN model’s input layer features and targets, the associated number of training epochs, and the corresponding learning rate strategy are detailed in Appendix B. In each instance, both the indicative numerical code and the optimal parameters $\theta^{\star}$ associated with the trained model will be accessible in the following GitHub repository: https://github.com/gregoryPapi/deep-universal-relations.git.
@@ IV.1.1 Relation for the star’s normalized moment of Inertia
Above all, the star’s moment of inertia strongly depends on the EoS. In the case of slow rotation, they are well-established EoS-insensitive relations for the reduced $\bar{I}$ with accuracy $\mathcal{O}(1\%)$ [149, 120, 132, 137, 104]. In addition, significant efforts have also been devoted to estimating this global parameter for rapidly rotating NSs (see, for example, the Refs. [143, 141, 142, 153] for a review). However, in the case of rapid rotation, there is still a significant demand for a highly accurate universal formula to infer the stellar object’s moment of inertia. This formula must accommodate a broad range of EoS candidates, ensuring compatibility with multimessenger astronomy data, and encompasses various microphysical structures. Therefore, for our extended ensemble of realistic EoSs, it is essential to establish an accurate and effective universal relation for $\bar{I}$, encompassing a wide range of rotation rates, incorporating the rapid rotation case.
@@
The evaluation measures for the optimal model’s weights $\theta^{\star}$ associated by Eq. (11) on the training and validation sets are presented in Table (3).
@@
Figure 3: Absolute relative error distribution ($100\%\times({\bar{I}}_{\mathrm{model}}-\bar{I})/\bar{I})$ derived evaluating the suggested regression model (11) on the test set. The error variance of the model’s predictions is displayed on a logarithmic scale. In addition, relative deviations for fitting functions proposed in the literature are shown in different colors, providing a basis for comparative analysis.
In the constructed model’s wireframe representation [^7] To visualize the ANN-predicted relation between the input parameters and the target quantity, we constructed a wireframe surface as follows: a uniform mesh grid was generated over the 2D input space ($\chi,\bar{Q}$), using the mesh_grid_funct in python, which samples values within the domain of the test dataset. For each grid point, the trained neural network predicted the corresponding target value (e.g., $\bar{I}$), producing a 2D array of model estimates via the ANN model function. The resulting surface was plotted as a semi-transparent wireframe over the 3D scatter plot of the actual test data, allowing for direct visual comparison between the model predictions and the ground truth., the test data lies approximately on a single universal plane, while the varying colors of the depicted data correspond to different values of the star’s compactness $C$, as indicated in the accompanying vertical colored bar. As anticipated, stellar modes that are less compact, corresponding to higher values of the reduced quadrupole deformation, demonstrate larger normalized moments of inertia. Additionally, the predicted wireframe includes regions that lie outside the physically meaningful parameter space. When the model is evaluated on inputs beyond the domain of the test dataset, it produces a smooth extrapolation of the surface—despite the fact that no physically viable stellar models exist in that regime. This behavior does not indicate a flaw in the model, but rather corroborates the robustness of the learned representation.
@@
Building upon this, we employed the trained DL model (11) to interpret its predictions on a test subset $X_{1}$ consisting of $10000$ NS configurations. Given the computational demands of the SHAP procedure [154, 155], this data sample was randomly selected [^8] As before, the random.seed(42) function in Python was applied to ensure reproducibility across experimental runs. from the initial test dataset to expedite the analysis. We have to note that the selected dataset encompasses values of $\chi\in[0.141,0.774]$ and $\bar{Q}\in[1.343,14.370]$, ensuring adequate representation of the parameter space. Subsequently, we compute the SHAP values for this subset to interpret the contribution of each input feature to the model’s predictions. Fig. 5 presents the SHAP values, as defined in Eq. (4), illustrating the influence of each parameter on the predictions for every stellar model in the examined subset.
Figure 5: SHAP analysis, based on Eq. (4), illustrating the contribution of each feature to the predictions of the trained model (11) for the selected subset of NS configurations. The color scale indicates the magnitude of each feature’s value. Top panel: Contribution of each feature to the model’s prediction results. The horizontal axis is given in logarithmic scale. Bottom panel: Feature importance based on average SHAP values indicating the parameters’ average impact on model predictions.
@@
Furthermore, as shown in the bottom panel, we present the mean absolute SHAP values for each feature, averaged across the NS models in the test subset. The rotation parameter $\chi$ corresponds to a mean value of $0.78$, whereas the quadrupole deformation parameter $\bar{Q}$ attains a substantially higher value of $4.32$. These findings indicate that both parameters play a role in guiding the model’s predictive performance. Specifically, $\chi$ characterizes the rotational state of the star, enabling the network to capture rotational patterns within the data while $\bar{Q}$ encapsulates the deformation induced by rotation, thereby providing complementary structural information. Notably, the influence of $\bar{Q}$ is closely tied to $\chi$, as $\bar{Q}$ not only reflects rotational deformation but also captures sensitivity to the underlying EoS. This dual dependence renders $\bar{Q}$ as a particularly informative parameter for pattern recognition, especially in the context of slowly rotating NSs where established EoS-independence between $\bar{Q}$ and $\bar{I}$ become prominent. The increased mean SHAP value attributed to $\bar{Q}$ underscores its enhanced role in refining predictive accuracy, particularly in regimes of higher rotation where deformation effects are more significant. In any case, the interplay between $\chi$ and $\bar{Q}$ strengthens the model’s predictive capability, ensuring robust universal inference across a diverse range of stellar configurations.

### v2

@@ Abstract
Relations between stellar properties independent of the nuclear equation of state offer profound insights into neutron star physics and have practical applications in data analysis by breaking the degeneracy between the parameters of interest. Commonly, these relations are derived from utilizing various realistic nuclear cold hadronic, hyperonic, and hybrid EoS models, each of which should obey the current multimessenger constraints and cover a wide range of stiffnesses. Concurrently, the field of multimessenger astronomy has been significantly enhanced by the advent of gravitational wave astronomy, which increasingly incorporates deep learning techniques and algorithms. At the same time, X-ray spectral data from NICER based on known pulsars are available, and additional observations are expected from upcoming missions. Therefore, in the coming years, a wealth of information on neutron stars is expected to be gathered from multiple observational channels. In this study, we revisit established universal relations, introduce new ones, and reassess them using a feed-forward neural network as a regression model. More specifically, we mainly propose “deep” EoS-insensitive hypersurface relations for rapidly rotating compact objects between several of the star’s global parameters, such as the moment of inertia, the reciprocal of the moment of inertia, the rotational frequency, mass and current multipole moments and the equatorial radius, which achieve an accuracy of within $1\%$ in most cases, with only a small fraction of investigated models exceeding this threshold. While analytical expressions can be used to represent some of these relations, the neural network approach demonstrates superior performance, particularly in complex regions of the parameter space. Furthermore, we use the SHapley Additive exPlanations (SHAP) method to interpret the predictions of the suggested feed-forward network, as SHAP is grounded in a robust theoretical framework inspired by the field of cooperative Game Theory. Most importantly, these highly accurate universal relations, empowered with the interpretability description, could be used in efforts to constrain the high-density equation of state in neutron stars, with the potential to enhance our understanding as new observables emerge.
@@ III.1 Mapping explainability using the SHAP method
Unfortunately, recent ML problems are becoming increasingly challenging and require more advanced and complex models to tackle them [184]. However, model complexity typically presents a trade-off between explainability and performance in decision-making. For instance, simpler models like those employing linear regression are highly interpretable but typically fall short in predictive performance compared to DL ones. In addition, DL models often sacrifice interpretability due to their inherent complexity, and nonlinear transformations, among other factors [185]. On top of that, contrary to the common perception that DL models operate as “black-boxes”, they are, in fact, well-defined systems, and we can provide clear explanations for how the input parameters influence their behavior. It should be noted that in several cases, understanding why a model made a certain prediction can be crucial.Over the years, various techniques have been developed to interpret the predictions of complex models, including LIME, DeepLIFT, and SHAP [154]. Among these, SHapley Additive exPlanations (SHAP) is a widely recognized method inspired by the concept of Shapley values from the field of cooperative game theory [155, 156].
Building on this foundation, it is essential to demonstrate how Shapley values, denoted as $\phi_{i}(x)$, provide a systematic quantification of the contribution of each feature $f_{i}$ to the model’s prediction for a given input $x$. At first, let us consider a set of features $D=\{f_{1},\dots,f_{d}\}$ with dimension $|D|=d$. For each subset of features $S\subseteq D$, we can define a model $\mathcal{F}_{S}:\mathbb{R}^{|S|}\rightarrow\mathbb{R}$, which only takes the values of the features in $S$. Building on this framework, we define
$\mathcal{F}:\bigcup_{S\subseteq[d]}(S\times\mathbb{R}^{|S|})\to\mathbb{R},$ (4)
where $\mathcal{F}(\emptyset)=0$ and $d]=\{1,\dots,d\}$. The Shapley value for feature $f_{i}$ is then defined as:
$\phi_{i}(x)=$
$\sum_{S\subseteq[d]\setminus\{f_{i}\}}\frac{|S|!(d-|S|-1)!}{d!}\Big(\mathcal{F}_{S\cup\{f_{i}\}}(x_{S\cup\{f_{i}\}})-\mathcal{F}_{S}(x_{S})\Big)$ (5)
where the summation iterates over all subsets $S$ that do not contain $f_{i}$. Note that $x_{S}$ denotes the values of the features in $S$, and $\mathcal{F}_{S}(x_{S})$ is the corresponding model output. Therefore, this computes the average marginal contribution of feature $f_{i}$ across all possible subsets. This approach provides a rigorous framework for assessing the contribution of each feature to the ML model’s predictions.
@@
$\hat{\mathcal{F}}(x)=\mathbb{E}_{X}(\hat{\mathcal{F}}(X))+\sum_{j=1}^{d}\phi_{j}(x),$ (6)
where $\hat{\mathcal{F}}(x)$ represents the model’s prediction, $\mathbb{E}_{X}(\hat{\mathcal{F}}(X))$ is the model’s expected value, and $\phi_{j}$ [^3] In general, SHAP value $\phi_{i}(x)$ is a function of the model’s feature vector $x$. Nevertheless, for simplicity, we adopt the more straightforward notation $\phi$ as a definition. are the individual SHAP values. In Fig. 1, as an indicative clarification example, we present the contribution of each SHAP value $\phi_{i}$ regarding the specific input vector $x_{i}=(f_{i,1},f_{i,2},f_{i,3},f_{i,4})$, containing four features. Each SHAP value contributes positively (represented by a green-colored vector) or negatively (represented by a red-colored vector) to the model’s inference, $\hat{\mathcal{F}}(x_{i})$, relatively to $\mathbb{E}_{X}(\hat{\mathcal{F}}(X))$. These values serve as useful tools to explain the model’s prediction ability, offering interpretability to the deep learning model’s predictions.
@@
In this work, SHAP is applied in the framework of NS universal relations, where physical quantities associated with the star’s stellar global parameters are used as feature inputs. The resulting SHAP value attributions are defined for a specific trained DL model and its corresponding feature set. Therefore, SHAP values are model-dependent interpretability scores, defined with respect to a specific trained model. For instance, retraining can lead to different learned optimal $\theta^{\star}$ parameters (e.g., due to stochastic optimization or convergence to a different local optimum), and consequently to some changes in the SHAP attributions across features. Having a well-trained model, it is important to note that SHAP values are not measures of the model’s accuracy; predictive performance should instead be assessed separately using standard statistical evaluation measures. Depending on the parametrization, one or more features may dominate the average attribution, or several may contribute comparably. Comparable SHAP values across features do not imply higher or lower model’s inference accuracy; rather, they reflect how the associated optimal weights distribute sensitivity across inputs. In any case, inference accuracy ultimately depends on the quality of the training data, the suitability of the parametrization, and the characteristics of the model chosen to verify the data. Therefore, SHAP is not intended to assess model performance, but rather to interpret how trained DL models distribute importance across input features.
In summary, interpreting the results of a DL model using SHAP values provides valuable insights into the factors influencing its predictions. SHAP values offer a systematic approach to quantifying the contribution of each input feature to the model’s output, thereby aiding in understanding its decision-making process. By analyzing SHAP values, we can identify which features have the most significant impact (in terms of magnitude) and how they contribute (positively or negatively) relative to the model’s expected value. This interpretability framework not only enhances our understanding of the model’s behavior but also supports its validation and helps identify potential areas for improvement or further investigation. For instance, if a feature has a very small mean SHAP value, one might consider excluding that input and retraining the model with a simpler parametrization. Such steps, however, should be approached with caution. Features are fundamental to model inference, and an EoS-insensitive parametrization allows the DL model to capture patterns in the data and achieve reliable performance. Removing a feature can alter the optimization landscape, potentially making training less stable and reducing the model’s predictive accuracy or generalization. Therefore, any feature exclusion should be carefully validated by benchmarking against standard evaluation metrics, ideally using cross-validation.
@@ III.2 Deep Learning Architecture, Training and Testing
$\mathcal{L}_{\delta}\left(z,\hat{z}(\theta)\right)=\sum_{i=1}^{n}\mathcal{L}_{i},$ (7)
where each term $\mathcal{L}_{i}$ is given by:
$\mathcal{L}_{i}=\begin{cases}\frac{1}{2}|z_{i}-\hat{z}_{i}(\theta)|^{2}&\text{, }|z_{i}-\hat{z}_{i}|\leq\delta,\\ \\ \delta\left(|z_{i}-\hat{z}_{i}(\theta)|-\frac{1}{2}\delta\right)&\text{, }|z_{i}-\hat{z}_{i}|>\delta,\end{cases}$ (8)
@@
For designing a DL architecture that satisfies our requirements, we utilized the PyTorch library [189]. For each EoS of cold, dense nuclear matter included in our catalog, we have used a random selection of 25% rotating stellar models as a test set444random.seed(42) method implemented in Python was used to reproduce the same sample of random NS configurations for testing. This particular choice ensures that the results are consistent across different runs.. In our systematic investigation, we employed an ANN architecture described by an input layer encompassing the features [^5] Depending the case, the specific characteristics of these features will be clarified in the Sec. IV. $x=(f_{1},...,f_{d})$, followed by five hidden layers denoted as $H_{1},\ldots,H_{5}$, while concluding in a singular output layer represented by $\hat{z}$. The number of the associated neurons regarding each hidden layer is shown in Table (1).
@@
to map each input feature $f_{i}$ employed to the interval $[0,1]$. This process is carried out on the entire dataset, which includes the associated quantities. Feature scaling is essential to ensure that all features contribute equally, preventing the dominance of those with larger values. This practice enhances the convergence and overall performance of the algorithms used. Either way, this process is an important step to render the learning process faster. Depending on the universal relation, we employ an adaptive learning rate, a tailored learning rate strategy, a case-favored mini-batch size, and a specific number of epochs for each training phase. In each case, the training time is linearly proportional to the number of epochs.
Across all of the regression models examined, we employed comprehensive evaluation measures to assess the performance of each ANN across the associated test sets (see e.g., [143, 153] for a detailed review). The statistical measures utilized include Mean Absolute Error (MAE), Max Residual Error (MRE), Mean Squared Error (MSE), Max Relative Deviation ($\mathrm{d}_{\text{max}}$), Mean Absolute Percentage Error (MAPE), Explained Variance, and Determination Coefficient ($R^{2}$) [^6] For more details about the statistical evaluation functions the interested reader can be directed to this link: https://scikit-learn.org/stable/modules/model_evaluation.html. Notably, the evaluation measures in the validation set show a slight degradation in performance compared to the training set, as expected. In each case, as is standard practice, we report the evaluation measures based on the performance observed on the test datasets.
@@
Furthermore, Fig. 2 shows the $\chi-C$, $\chi-\log\bar{I}$, $\chi-\log\bar{Q}$, and $\chi-\log\bar{S}_{3}$ representations for each EoS, based on an indicative subsample of $40000$ NS configurations from the full dataset. For visualization purposes, this subsample corresponds to a more distinct set of polar-to-equatorial radius ratios $r_{\mathrm{pole}}/r_{\mathrm{eq}}$ than that employed in the full dataset.
Figure 2: Distributions of the $\chi-C$, $\chi-\log\bar{I}$, $\chi-\log\bar{Q}$, and $\chi-\log\bar{S}_{3}$ parameter spaces, encompassing a wide range of rotation rates and degree of stiffness. Each color represents the EOS mapping as highlighted in Fig. 31 of Appendix A.
@@
Last but not least, we define a relation between the NS parameters as EoS-insensitive when the selected features at the DL model’s inference yield relative errors in the training and test sets of $\lesssim\mathcal{O}(10\%)$. Crucially, our regression “deep” models accurately reproduce the data on the test set, showcasing a state-of-the-art generalization capability beyond the training data. For each universal relation examined, the ANN model’s input layer features and targets, the associated number of training epochs, and the corresponding learning rate strategy are detailed in Appendix B. In each instance, both the indicative numerical code and the optimal parameters $\theta^{\star}$ associated with the trained model are publicly available in Ref. [193].
@@ IV.1.1 Relation for the star’s normalized moment of Inertia
Above all, the star’s moment of inertia strongly depends on the EoS. In the case of slow rotation, there are well-established EoS-insensitive relations for the reduced $\bar{I}$ with accuracy $\mathcal{O}(1\%)$ [149, 120, 132, 137, 104]. In addition, significant efforts have also been devoted to estimating this global parameter for rapidly rotating NSs (see, for example, the Refs. [143, 141, 142, 153] for a review). However, in the case of rapid rotation, there is still a significant demand for a highly accurate universal formula to infer the stellar object’s moment of inertia. This formula must accommodate a broad range of EoS candidates, ensuring compatibility with multimessenger astronomy data, and encompasses various microphysical structures. Therefore, for our extended ensemble of realistic EoSs, it is essential to establish an accurate and effective universal relation for $\bar{I}$, encompassing a wide range of rotation rates, incorporating the rapid rotation case.
@@
The evaluation measures for the optimal model’s weights $\theta^{\star}$ associated with Eq. (13) on the validation set are presented in Table (3).
@@
Figure 4: Absolute relative error distribution ($100\%\times({\bar{I}}_{\mathrm{model}}-\bar{I})/\bar{I})$ derived evaluating the suggested regression model (13) on the test set. The vertical axis corresponds to the probability density function (PDF) of the test dataset, shown on a logarithmic scale. In addition, relative deviations for fitting functions proposed in the literature are shown in different colors, providing a basis for comparative analysis.
In the constructed model’s wireframe representation [^8] To visualize the ANN-predicted relation between the input parameters and the target quantity, we constructed a wireframe surface as follows: a uniform mesh grid was generated over the 2D input space ($\chi,\bar{Q}$), using the mesh_grid_funct in python, which samples values within the domain of the test dataset. For each grid point, the trained neural network predicted the corresponding target value (e.g., $\bar{I}$), producing a 2D array of model estimates via the ANN model function. The resulting surface was plotted as a semi-transparent wireframe over the 3D scatter plot of the actual test data, allowing for direct visual comparison between the model predictions and the ground truth., the test data lies approximately on a single universal plane, while the varying colors of the depicted data correspond to different values of the star’s compactness $C$, as indicated in the accompanying vertical colored bar. As anticipated, stellar modes that are less compact, corresponding to higher values of the reduced quadrupole deformation, demonstrate larger normalized moments of inertia. Additionally, the predicted wireframe includes regions that lie outside the physically meaningful parameter space. When the model is evaluated on inputs beyond the domain of the test dataset, it produces a smooth extrapolation of the surface—despite the fact that no physically viable stellar models exist in that regime. This behavior does not indicate a flaw in the model but simply reflects the smooth continuation of the learned mapping beyond the physical domain, where no viable stellar models exist.
@@
Building upon this, we employed the trained DL model (13) to interpret its predictions on a test subset $X_{1}$ consisting of $10000$ NS configurations. Given the computational demands of the SHAP procedure [154, 155], this data sample was randomly selected [^9] As before, the random.seed(42) function in Python was applied to ensure reproducibility across experimental runs. from the initial test dataset to expedite the analysis. We have to note that the selected dataset encompasses values of $\chi\in[0.141,0.774]$ and $\bar{Q}\in[1.343,14.370]$, ensuring adequate representation of the parameter space. Subsequently, we compute the SHAP values for this subset to interpret the contribution of each input feature to the model’s predictions. Fig. 6 presents the SHAP values, as defined in Eq. (III.1), illustrating the influence of each parameter on the predictions for every stellar model in the examined subset.
Figure 6: SHAP analysis, based on Eq. (III.1), illustrating the contribution of each feature to the predictions of the trained model (13) for the selected subset of NS configurations. The color scale indicates the magnitude of each feature’s value. Top panel: Contribution of each feature to the model’s prediction results. Bottom panel: Feature importance based on average SHAP values indicating the parameters’ average impact on model predictions.
@@
Furthermore, as shown in the bottom panel, we present the mean absolute SHAP values for each feature, averaged across the NS models in the test subset. The rotation parameter $\chi$ corresponds to a mean value of $0.78$, whereas the quadrupole deformation parameter $\bar{Q}$ attains a substantially higher value of $4.32$. These findings indicate that both parameters play a role in exerting a measurable influence on the model’s predictive performance. Specifically, $\chi$ characterizes the rotational state of the star, enabling the network to capture rotational patterns within the data while $\bar{Q}$ encapsulates the deformation induced by rotation, thereby providing complementary structural information. This dual dependence renders $\bar{Q}$ as a particularly informative parameter for pattern recognition—especially in the slow-rotation limit and along NS sequences at fixed $\chi$, where the established EoS-independence between $\bar{Q}$ and $\bar{I}$ becomes prominent. The increased mean SHAP value attributed to $\bar{Q}$ indicates that the trained model relies more strongly on this parameter when forming its predictions, particularly in regimes of higher rotation where deformation effects are more significant. In any case, the interplay between $\chi$ and $\bar{Q}$ strengthens the model’s predictive capability, ensuring robust universal inference across a diverse range of stellar configurations.